Math 46 Homework Assignment (11/16 – 11/23) Assignment will be due on 11/30 the Tuesday during the lecture.

Key notions introduced this week: (1) Variations of Parameters;

(2) Wronskian;

(3) Determinants;

(4) Initial value problems;

(5) Solution of system of two linear equations.

Homework problems:

Chapter 12. 11, 13, 19, 23. Chapter 13. 7, 12, 15.